Calculate the number of moles of methyl alcohol in 200ml substance. (Density = 0.2715 gram /
cm3
)
1
Expert's answer
2020-09-23T04:39:37-0400

find the mass of alcohol m=V *d=200*0.2715=53.2 g

find Mr(CH3OH)= Ar(C)+4Ar(H)+Ar(O)=12+4+16=32

find amount of moles n= m/Mr=53.2/32=1.7 mol